OpenStudy (anonymous):

Which ordered pair is a solution of the equation y = –9x + 4? (10, –86) (–4, –58) (6, –41) (–6, 57)

OpenStudy (anonymous):

(10,86) y=-9(10)+4 = -90+4 y=-86

OpenStudy (anonymous):

sorry, it's (10,-86)